OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x – 2y = 2 x + 3y = 17 a (2y + 2) + 3y = 17 b (–2y + 2) + 3y = 17 c x + 3(2y + 2) = 17 d x + 3(–2y + 2) = 17

OpenStudy (anonymous):

a

OpenStudy (vishweshshrimali5):

a

OpenStudy (vishweshshrimali5):

a a (2y + 2) + 3y = 17

OpenStudy (vishweshshrimali5):

a a (2y + 2) + 3y = 17

OpenStudy (vishweshshrimali5):

a a (2y + 2) + 3y = 17

OpenStudy (vishweshshrimali5):

a a (2y + 2) + 3y = 17

OpenStudy (vishweshshrimali5):

a a (2y + 2) + 3y = 17

OpenStudy (vishweshshrimali5):

a a (2y + 2) + 3y = 17

OpenStudy (vishweshshrimali5):

a a (2y + 2) + 3y = 17

OpenStudy (anonymous):

Thanks